the length of a rectangular garden is 4 yards more than its width. The area of the garden is 60 square yards. Find the dimensions of the garden.

Answer by TimothyLamb(4379)

L = W + 4
LW = 60
---
LW = 60
(W + 4)W = 60
W^2 + 4W + (-60) = 0
---
above is in quadratic equation form: ax^2 + bx + c = 0
---
use the quadratic formula to solve for two solutions:
x = ( -b +- sqrt( bb - 4ac ) ) / 2a
---
the two solutions are:
W = 6
W = -10
the negative solution doesn't make sense for length, so use the positive solution.
---
L = 10
W = 6
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php